A bottle of a good wine originally cost $36. For each of the following scenarios, write a formula for the value in dollars W of the bottle of wine as a function of the number of years t.

(a)

The price of the bottle increased by 13% each year.

W =

(b)

The price of the bottle decreased by 7% each year.

W =

(c)

The price of the bottle dropped by $3 each year.

W =

(d)

The price of the bottle remained steady.

W =
Answer #1

a) Price of the bottle increased by 13% each year:

W = 36 * (1.13)t

b) Price of the bottle decreased by 7% each year:

W = 36 * (0.93)t

c) Price of the bottle dropped by $3 each year:

W = 36 - 3t

d) Price of the bottle remained steady:

W = 36